How many rectangular pieces of wood of size 20 cm by 45 cm can be cut from a plywood sheet of 120 cm by 240 cm?
A 29 B 30 C 31 D 32 E 33

Answer by Theo(13342)

cutting lengthwise, you can get 240 / 45 = 5 boards for each width.

The number of widths = 120 / 20 = 6

total boards = 5*6 = 30.

cutting crosswise, you can get 120 / 45 = 2 boards for each width.

the number of widths = 240 / 20 = 12

total boards = 12 * 2 = 24

since 30 is one of your selections and 24 is not, go with 30.

answer is selection B.